Calculate the amount of chlorine gas evolved during the electrolysis of molten potassium

chloride on passing a current of 0.5 A for two hours through it.


1
Expert's answer
2021-01-18T01:34:53-0500

n = I*t/F/z = 0.5*2*60*60/96500/2 = 0.01865 mol